OpenStudy (anonymous):

Find Cartesian equation for a curve that contains the parametrized curve. x = 4sin t, y = 2cos t 0≤t≤π

OpenStudy (anonymous):

x = 4sin t y = 2cost x/4 = sint --->1 y/2 = cost --->2 square equation 1 and 2 (x/4)^2 = (sint)^2 --->3 (y/2)^2 = (cost)^2 ---->4 add eq 3 and 4 (x/4)^2 +(y/2)^2 = (sint)^2 + (cost)^2 as (sint)^2 + (cost)^2 = 1 thus (x/4)^2 +(y/2)^2 = 1

OpenStudy (anonymous):

Omg this seems so easy now, I feel so dumb thank you life saver.

OpenStudy (anonymous):

you are welcomeeeeeeeeeeeeeeeee